# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 7 Conclusion
Rare event sampling is a promising strategy to study extreme weather more efficiently with computer models by repeatedly cloning, perturbing, and re-simulating the most extreme events in an ensemble while tracking statistical weights. However, sudden and transient events such as mid-latitude precipitation present a particular challenge for rare event algorithms, leaving ensembles little time to diversify before the event passes by. Ensemble boosting (Gessner et al.,, 2021; Gessner,, 2022; Fischer et al.,, 2023; Bloin-Wibe et al.,, 2025) and “trying-early adaptive multilevel splitting” (TEAMS; Finkel and O’Gorman,, 2024) get around this problem by perturbing events farther in advance by some advance split time (AST) to allow ensembles to spread, but this opens a pivotal question: how should we choose the AST for maximal accuracy and efficiency? If AST is too short, perturbations can’t grow enough to give useful samples, and if it is too long, they regress to climatology. To deploy advance-splitting methods at scale, we need more reliable ways to set the AST as well as other hyperparameters.
In this paper, we have established the conditionally optimal advance split time (COAST) as an intrinsic quantity, not to the whimsies of a particular algorithm but to the dynamical system itself, as well as the target observable of interest, the imposed distribution over perturbations, and the initial conditions which may vary in their predictability. We formulate COAST mathematically as the solution an optimization problem, and through a systematic boosting-based sampling and estimation procedure we discern the optimization landscape in the context of an idealized physical model: a baroclinically unstable quasi-geostrophic flow, with local passive tracer fluctuations as our extreme event of interest. To faciliatate more efficient rare event sampling applications, we have further proposed various parsimonious rules for finding the COAST, and evaluated these rules empirically in the QG model.
We have three conclusions to report, one physical and two algorithmic:
@@
    An optimal AST exists and is strictly between zero and infinity, consistently across many target locations in the channel domain. It varies slowly with latitude, appearing (smaller, larger) in regions of (negative, positive) meridional wind shear, e.g., the (northern, southern) edges of westerly jets.
@@
    Several different rules for selecting the COAST are equally effective. Beyond the simplest option of setting a single fixed AST (called $A^{\$}$), one can set a conditional AST (called $A^{\text{\textcent}}$) by thresholding on ensemble dispersion. Both $A^{\$}$ and $A^{\text{\textcent}}$ perform similarly at tail reconstruction, but both unfortunately require an arbitrary threshold choice, which there is no established method for selecting. Here we selected thresholds post hoc with knowledge of the ground truth. The tentative rule proposed in Finkel and O’Gorman, (2024)—that $A^{\$}\approx$ the time until ensembles disperse to $\frac{3}{8}$ their saturation value—appears to be the best possible single choice, but further improvement is possible by tailoring AST to the target location and the initial condition.
@@
There are many important avenues of research indicated by the present study, both methodology-oriented and science-oriented. On the algorithmic front, it remains to be seen whether thresholded entropy succeeds at matching tail statistics in general systems, but the consistency across different targets within the QG model is encouraging. We suspect that some objective function over distributions is broadly applicable. Furthermore, the shape of perturbations is a possibly very important lever on the potency of perturbations, acting in concert with their timing. While we limited our present study to a two-dimensional perturbation space based on linearized dynamics about a state of rest, a natural extension would be to use flow-dependent singular vectors as in operational weather forecasting. By design, they effect faster ensemble spread in the small-perturbation regime; however, it must be checked if their advantages carry into the finite-amplitude regime needed for effective rare event sampling. Computational tools such as adjoints, especially in novel machine learning models, invite the use of gradient-based optimization (Wang et al.,, 2020; Vonich and Hakim,, 2024).
@@
Intriguing dynamical questions also arise from the latitude dependence of the COAST, which can be seen as a predictability index tailored to extremes: how do the physical parameters such as topography, rotation rate, and the spatial domain affect COAST? Is the effect entirely explainable through the extreme value statistics, as we have speculated, or can two similarly shaped tails belie extremely different COAST behavior? These questions merit further parameter exploration, both within and beyond the quasigeostrophic framework. We expect to draw insight from recent theoretical advances relating extreme value theory to the geometry of chaotic attractors (Lucarini et al.,, 2016).
In summary, our work makes empirical progress on important theoretical and algorithmic questions regarding the limits, and probabilities, of the most extreme weather events. We have established the existence of an optimization landscape, and only with this basic pre-requisite information can we proceed to efficiently optimize.
Code availability
@@
The code to generate all results is available at the Github COAST repository, specifically commit https://github.com/justinfocus12/COAST/commit/cda6c2c181739fc0f16cfc9d6b0d2369430e6e67. J.F. is happy to provide guidance on use and extension of the code.
Acknowledgements
@@
We thank Glenn Flierl, Andre Souza, and Talia Tamarin-Brodsky for helpful discussions and advice on theoretical and computational aspects of this work. This research is part of the MIT Climate Grand Challenge on Weather and Climate Extremes. Support was provided by Schmidt Sciences, LLC. Computations were performed on the MIT Engaging cluster.

### v2

@@ 7 Conclusion
Rare event sampling is a promising strategy to study extreme weather more efficiently with computer models by repeatedly cloning, perturbing, and re-simulating the most extreme events in an ensemble while tracking statistical weights. However, sudden and transient events such as mid-latitude precipitation present a particular challenge for rare event algorithms, leaving ensembles little time to diversify before the event passes by. Ensemble boosting (Gessner et al., 2021; Gessner, 2022; Fischer et al., 2023; Bloin-Wibe et al., 2025) and “trying-early adaptive multilevel splitting” (TEAMS; Finkel and O’Gorman, 2024) get around this problem by perturbing events farther in advance by some advance split time (AST) to allow ensembles to spread, but this opens a pivotal question: how should we choose the AST for maximal accuracy and efficiency? If AST is too short, perturbations can’t grow enough to give useful samples, and if it is too long, they regress to climatology. To deploy advance-splitting methods at scale, we need more reliable ways to set the AST as well as other hyperparameters.
In this paper, we have established the conditionally optimal advance split time (COAST) as a quantity more intrinsic to the dynamical system than to the whimsies of a particular rare event algorithm by removing the confounding effect of randomly selecting ensemble members to split. The COAST also depends on the target observable of interest, the imposed distribution over perturbations, and the initial conditions which may vary in their predictability. We formulate COAST mathematically as the solution an optimization problem, and through a systematic boosting-based sampling and estimation procedure we discern the optimization landscape in the context of an idealized physical model: a baroclinically unstable quasi-geostrophic flow, with local passive tracer fluctuations as our extreme event of interest. To faciliatate more efficient rare event sampling applications, we have further proposed various parsimonious rules for finding the COAST, and evaluated these rules empirically in the QG model.
We have four conclusions to report:
@@
    A boosting procedure, generated with a suitable AST, can well-approximate a probability distribution’s tail using MoCTail or PoPTail estimators.
@@
    The optimal AST is strictly greater than zero and varies slowly with latitude, appearing smaller in regions of negative meridional wind shear (e.g., the northern edges of westerly jets) and larger in regions of positive meridional wind shear (e.g., the southern edges of westerly jets).
@@
    Several different rules for selecting the COAST are equally effective. Beyond the simplest option of setting a single fixed AST (called $A^{\text{U}}$), one can set a conditional AST (called $A^{\text{PC}}$) by thresholding on ensemble dispersion. Both $A^{\text{U}}$ and $A^{\text{PC}}$ perform similarly at tail reconstruction, but both unfortunately require a threshold choice, which there is no established method for selecting. Here we selected thresholds post hoc with knowledge of the ground truth. The rule proposed in Finkel and O’Gorman (2024)—that $A^{\text{U}}\approx$ the time until ensembles disperse to $\frac{3}{8}$ their saturation value—appears to be a good single choice, but further improvement is possible by tailoring AST to the target location and the initial condition.
  4. 4.
@@
There are many important avenues of research indicated by the present study, both methodology-oriented and science-oriented. On the algorithmic front, it remains to be seen whether thresholded entropy succeeds at matching tail statistics in general systems, but the consistency across different targets within the QG model is encouraging. We suspect that some objective function over distributions is broadly applicable. Furthermore, the shape of perturbations is a possibly very important lever on the potency of perturbations, acting in concert with their timing. While we limited our present study to a two-dimensional perturbation space based on linearized dynamics about a baroclinically unstable background flow, a natural extension would be to use flow-dependent singular vectors as in operational weather forecasting. By design, they effect faster ensemble spread in the small-perturbation regime; however, it must be checked if their advantages carry into the finite-amplitude regime needed for effective rare event sampling. Computational tools such as adjoints, especially in novel machine learning models, invite the use of gradient-based optimization (Wang et al., 2020; Vonich and Hakim, 2024).
@@
Intriguing dynamical questions also arise from the latitude dependence of the COAST, which can be seen as a predictability index tailored to extremes: how do the physical parameters such as topography, rotation rate, and the spatial domain affect COAST? Is the effect entirely explainable through the extreme value statistics, as we have speculated, or can two similarly shaped tails belie extremely different COAST behavior? These questions merit further parameter exploration, both within and beyond the quasigeostrophic framework. We expect to draw insight from recent theoretical advances relating extreme value theory to the geometry of chaotic attractors (Lucarini et al., 2016).
In summary, our work makes empirical progress on important theoretical and algorithmic questions regarding the probabilities of the most extreme weather events. Perturbed ensemble forecasts of individual weather events are distinct from the climatological distribution, but here we have given quantitative evidence for a relationship between the two—so long as the perturbations are well-timed – that can be exploited for efficient risk analysis via judicious perturbed simulations. Our work has elucidated what it means to be “well-timed”, and furthermore provided quantitative optimization criteria for perturbation timing. Only with this basic pre-requisite information on what to optimize, should we proceed to invest effort into optimizing efficiently.
[…]
@@
The code to generate all results, including simulation, statistical analysis, and plotting, is available at the Zenodo repository COAST (justinfocus12, 2025), https://doi.org/10.5281/zenodo.17355215. J.F. is happy to provide guidance on use and extension of the code upon request.
Appendix A Langevin Model
The schematic in Fig.1 comes from Langevin dynamics, consisting of a single particle moving in one dimension with position $X(t)$ and momentum $Y(t)$ subject to a potential gradient force, friction, and stochastic Gaussian white-noise forcing $W(t)$:
$dX(t)$ $=\frac{1}{m}Y(t)\,dt$ (45)
$dY(t)$ $=\Big[-V^{\prime}(X(t))-\gamma Y(t)\Big]\,dt+\sigma\,dW(t)$ (46)
where the potential function $V(x)\text{ has a quadratic core and logarithmic wings,}$ (47)
$V(x)$ $=\begin{cases}\frac{\alpha+1}{\beta}\Big(\log(\epsilon)+\frac{(x/\epsilon)^{2}-1}{2}\Big)&|x|\leq\epsilon\\ \frac{\alpha+1}{\beta}\log|x|&|x|>\epsilon,\end{cases}$ (48)
which leads to a heavy-tailed (in $x$) steady-state probability density $p(x,y)\propto\exp\big[-\beta(V(x)+\frac{y^{2}}{2m})\big]\sim|x|^{-(\alpha+1)}$ for large $|x|$. Constant parameters are $\gamma=0.05$ for friction, $m=1.2$ for mass, $\sigma=0.005$ for stochastic forcing strength, $\epsilon=0.25$ for the extent of the quadratic core of the potential, $\alpha=3.1$ which sets the tail weight, and $\beta=2m\gamma/\sigma^{2}$ which is the inverse temperature.
[…]
Justin Finkel formulated the initial study, carried out numerical computations, and wrote the initial draft. Paul O’Gorman and Justin Finkel both contributed to refining the methodology and substantially revising the manuscript.
[…]
The authors declare no competing interests relevant to this study.
Acknowledgements.
@@
We thank Glenn Flierl, Andre Souza, and Talia Tamarin-Brodsky for helpful discussions and advice on theoretical and computational aspects of this work. This research is part of the MIT Climate Grand Challenge on Weather and Climate Extremes. Support was provided by Schmidt Sciences. Computations were performed on the MIT Engaging cluster.